OpenStudy (anonymous):

is the answer to this equation "3(x-1) =2x - y" y=-x+3?

OpenStudy (anonymous):

3(x-1) =2x - y 3x-3 = 2x - y 3x-2x-3 = -y y = -3x + 2x + 3 y = -x + 3 Yup! you got it!